What is the first three terms of n+3

Answer by ikleyn(52790)

.
What is the first three terms of n+3
--------------------------------------

n = 1 ---> n+3 = 1 + 3 = 4.

n = 2 ---> n+3 = 2 + 3 = 5.

n = 3 ---> n+3 = 3 + 3 = 6.